the volume of a can of pork and beans is directly proportional to the height of the can. If the volume of the can is 300 in^3. When its height is 10.62 in, find the volume of a can with height 15.92 in
Answer by rfer(16322)

300/10.62=28.25
28.25*15.92=450 in^3
-------------
10.62/300=15.92/x
10.62x=300*15.92
10.62x=4776
x=4776/10.62
x=450 in^3